# Topological regularity of Busemann spaces of nonpositive curvature

### v1

@@ 7.3. Applications
Next we prove Theorem 1.7. For the Gromov–Hausdorff convergence, we refer to [20, Chapter 7]. We also use some notions and results listed in [65, Section 4].
We first recall the following globalization, which is actually the final step of the proof of the Cartan–Hadamard theorem (see Section 2.3). For the convenience of the reader, we include the proof (cf. [79, Theorem 9.3.4]).
@@
Next we prove the homotopical stability of small metric spheres in GNPC spaces with respect to the Gromov–Hausdorff convergence. Compare with [65, Proposition 2.3(2)]. Strictly speaking, we need the notion of convex geodesic bicombing here, but we defer it to the next subsection.
@@
By Corollary 5.24 and Remark 5.4, $\partial B(p_{j},r)$ and $\partial B(p,r)$ are locally uniformly contractible in the sense of [65, Definition 4.1]. In addition, by the $\varepsilon$-openness of a strainer map, Proposition 5.17, $\partial B(p_{j},r)$ certainly converges to $\partial B(p,r)$ in the Gromov–Hausdorff topology (more precisely, any point of $\partial B(p,r)$ is a limit of points of $\partial B(p_{j},r)$). Therefore, by Petersen’s stability theorem [83, Theorem A] ([65, Theorem 4.2]), we conclude that $\partial B(p_{j},r)$ is homotopy equivalent to $\partial B(p,r)$, as desired.
@@
The above result is actually the special case of the homotopical stability of the fibers of strainer maps. See [64, Theorem 13.1] for more details.
Now we prove Theorem 1.7. The proof is almost the same as [65, Theorem 7.1, Corollary 7.2]. Recall that a G-space has injectivity radius $\geq r$ if every shortest path of length $<r$ extends uniquely to length $r$. In particular, every two points at distance $<r$ is joined by a unique shortest path.
@@
Since $X_{j}$ and $X$ are uniformly locally contractible in the sense of [65, Definition 4.1] and $X$ is finite-dimensional by Proposition 3.1, it immediately follows from Petersen’s stability theorem [83, Theorem A] that $X_{j}$ is homotopy equivalent to $X$ for any large $j$. More precisely, for any given $\varepsilon>0$, $X_{j}$ and $X$ are $\varepsilon$-equivalence, provided $j$ is large enough (see [65, Theorem 4.2]). In view of the $\alpha$-approximation theorem of Chapman–Ferry [27] (see also [65, Theorem 4.7] and references therein), it remains to show that $X$ is a topological manifold.
@@
In the above proof, we used the fact that $\partial B(p_{j},r_{p})$ and $\partial B(p,r_{p})$ are locally uniformly contractible in the sense of [65, Definition 4.1], which was shown by a strainer argument in the proof of Lemma 7.15. This fact also follows from the assumption that $X_{j}$ is a G-space with injectivity radius $\geq r>r_{p}$.
Indeed, by using Krakus’ retraction inequality [57] (see also [91, Proposition 2.3]), one can show that any point of $\partial B(p_{j},r_{p})$ has an arbitrarily small contractible neighborhood (see the proof of [91, Theorem 2.12]). Hence $\partial B(p_{j},r_{p})$ are locally uniformly contractible. Furthermore, by the stability of local uniform contractibility ([17, Section 16, Main Theorem]; see also [83, Section 5, Theorem], [42, Theorem 1.2]), we see that $\partial B(p,r_{p})$ is also locally uniformly contractible.
@@
Let $X\in\mathcal{G}(r,L)$. By Lemma 7.14 and the injectivity radius bound, every point of $X$ has a tiny ball of radius $r/100$. By the proof of Proposition 3.1 and the packing number bound, every tiny ball in $X$ of radius $r/100$ is $L$-doubling. In particular, the dimension of $X$ is bounded above by a constant depending only on $L$. Since $X$ is covered by at most $L$ tiny balls of radius $r/100$, this implies that for any given $\varepsilon>0$, the minimal number of $\varepsilon$-balls needed to cover $X$ is bounded above by a constant $N_{L}(\varepsilon)$ depending only on $L$ and $\varepsilon$. Therefore, the family $\mathcal{G}(r,L)$ is precompact in the Gromov–Hausdorff topology (cf. [64, Proposition 5.10]).
@@ 7.4. Generalization
On the other hand, the non-uniqueness of shortest paths typically results in the non-existence of a unique projection to a “$\sigma$-convex subset” (compare [79, Corollary 8.2.6, Proposition 8.4.8]). However, we have never used such a projection.
The advantage of dropping the uniqueness of shortest paths is that this weaker notion is preserved under the Gromov–Hausdorff convergence, which was essential in the previous subsection. Compare the following with [64, Example 4.3].
@@ 8. Problems
The first is the generalization of upper curvature bounds mentioned in Section 1.5. By modifying the BNPC condition, one can define Busemann spaces with curvature $\leq\kappa$ for arbitrary $\kappa\in\mathbb{R}$ (see Remark 2.5). For example, any reversible Finsler manifold with flag curvature $\leq\kappa$ is a Busemann space with curvature $\leq\kappa$ locally ([55]).
@@
The next three problems are all related to the tangent cones of GNPC spaces, which also generalize previous results for GCBA spaces by Lytchak–Nagano [64]. Before discussing the problems, we briefly review the original results in [64] (cf. [78], [76], [15]). For analogous results for Alexandrov spaces with curvature bounded below, see [77], [75], [76], [82], [13].
@@
We say that $p$ is a $k$-regular point if $T_{p}X$ is isometric to $\mathbb{R}^{k}$ (with Euclidean norm). The set $R^{k}$ of $k$-regular points in $X$ is dense and has full measure with respect to the $k$-dimensional Hausdorff measure in the $k$-dimensional part $X^{k}$ of $X$. Here, the $k$-dimensional part $X^{k}$ is the set of points in $X$ with $\dim T_{p}X=k$ (in the sense of the Hausdorff or topological dimension).
@@
We want to develop GNPC analogs of the above results. For simplicity, we ignore the dimensional inhomogeneity of a GNPC space and formulate rough problems. The starting point would be the structure of the tangent cone.
@@
Describe the structure of the tangent cone of a GNPC space. Define regular points of a GNPC space.
To define regular points, we would need some standard norm, which plays the same role as the Euclidean norm for GCBA spaces. In general, the norm of the tangent space of a Finsler manifold depends on each point. However, if a Finsler manifold satisfies the BNPC condition, then the Finsler metric belongs to a special class called a Berward metric, as shown by Ivanov–Lytchak [51]. One of the properties of a Berwald metric is that the norm of the tangent space is all isometric. From our viewpoint that GNPC spaces is a generalization of Finsler manifolds with nonpositive flag curvature, it is natural to ask the following.
@@
Does a GNPC space satisfy any kind of Berwardness? For example, is the tangent cone isometric to a fixed (strictly convex) normed space for almost every point?
Related problems and partial counterexamples can be found in [51, Sections 1, 6]. Note that, unlike Riemannian manifolds, Finsler manifolds do not generally have a “canonical” measure, but for (reversible) Berwald spaces the Hausdorff measure can be regarded as a canonical measure (see [74, Section 10.2]). Therefore it would be natural to consider the Hausdorff measure for our GNPC spaces.
A more ambitious (or optimistic) problem would be the following. Compare with Remark 5.31.
@@
Does a GNPC space admit any kind of Finsler structure almost everywhere? Does a regular point have a neighborhood almost isometric to an open subset of the “standard” normed space?
@@
Another problem we would like to mention here is the diffeomorphism version of Theorem 1.3, which is indeed open even in the CAT(00) case. See [30] and [89] for related results.
@@
Let $X$ be a smooth $4$-manifold with a CAT(00)/BNPC metric. Is $X$ diffeomorphic to the standard $\mathbb{R}^{4}$?
Finally, we discuss Busemann spaces with lower curvature bounds. Recall that Theorems 1.1 and 1.2 were originally proved by Wu [96] for Alexandrov spaces with curvature bounded below. Reversing the inequality for comparison angles in Remark 2.5, one can also define Busemann spaces with curvature $\geq\kappa$, where $\kappa\in\mathbb{R}$ (cf. [45]). Clearly, every Alexandrov space is a Busemann space with the same lower curvature bound. Therefore it would be natural to ask
@@
Note that Wu’s proof [96] (cf. [97]) relies heavily on the structure theorem for Alexandrov spaces by Perelman [80], [81], which seems very hard to generalize to Busemann spaces. We point out that Lytchak–Nagano’s proof [65] can also be applied to Alexandrov spaces, by using the gradient flow of the distance function (see [84]) instead of the geodesic contraction we frequently used. However, to the best of our knowledge, it is not even known whether a Busemann space with a lower curvature bound is locally contractible. We remark that there is work of Kell [53] on Busemann spaces with nonnegative curvature. See also [54] for such Finsler manifolds.

### v2

@@ 7.3. Applications
Next we prove Theorem 1.7. For the Gromov–Hausdorff convergence, we refer to [20, Chapter 7]. We also use some notions and results listed in [64, Section 4].
We first recall the following globalization, which is actually the final step of the proof of the Cartan–Hadamard theorem (see Section 2.3). For the convenience of the reader, we include the proof (cf. [78, Theorem 9.3.4]).
@@
Next we prove the homotopical stability of small metric spheres in GNPC spaces with respect to the Gromov–Hausdorff convergence. Compare with [64, Proposition 2.3(2)]. Strictly speaking, we need the notion of convex geodesic bicombing here, but we defer it to the next subsection.
@@
By Corollary 5.24 and Remark 5.4, $\partial B(p_{j},r)$ and $\partial B(p,r)$ are locally uniformly contractible in the sense of [64, Definition 4.1]. In addition, by the $\varepsilon$-openness of a strainer map, Proposition 5.17, $\partial B(p_{j},r)$ certainly converges to $\partial B(p,r)$ in the Gromov–Hausdorff topology (more precisely, any point of $\partial B(p,r)$ is a limit of points of $\partial B(p_{j},r)$). Therefore, by Petersen’s stability theorem [82, Theorem A] ([64, Theorem 4.2]), we conclude that $\partial B(p_{j},r)$ is homotopy equivalent to $\partial B(p,r)$, as desired.
@@
The above result is actually the special case of the homotopical stability of the fibers of strainer maps. See [63, Theorem 13.1] for more details.
Now we prove Theorem 1.7. The proof is almost the same as [64, Theorem 7.1, Corollary 7.2]. Recall that a G-space has injectivity radius $\geq r$ if every shortest path of length $<r$ extends uniquely to length $r$. In particular, every two points at distance $<r$ is joined by a unique shortest path.
@@
Since $X_{j}$ and $X$ are uniformly locally contractible in the sense of [64, Definition 4.1] and $X$ is finite-dimensional by Proposition 3.1, it immediately follows from Petersen’s stability theorem [82, Theorem A] that $X_{j}$ is homotopy equivalent to $X$ for any large $j$. More precisely, for any given $\varepsilon>0$, $X_{j}$ and $X$ are $\varepsilon$-equivalence, provided $j$ is large enough (see [64, Theorem 4.2]). In view of the $\alpha$-approximation theorem of Chapman–Ferry [27] (see also [64, Theorem 4.7] and references therein), it remains to show that $X$ is a topological manifold.
@@
In the above proof, we used the fact that $\partial B(p_{j},r_{p})$ and $\partial B(p,r_{p})$ are locally uniformly contractible in the sense of [64, Definition 4.1], which was shown by a strainer argument in the proof of Lemma 7.15. This fact also follows from the assumption that $X_{j}$ is a G-space with injectivity radius $\geq r>r_{p}$.
Indeed, by using Krakus’ retraction inequality [55] (see also [91, Proposition 2.3]), one can show that any point of $\partial B(p_{j},r_{p})$ has an arbitrarily small contractible neighborhood (see the proof of [91, Theorem 2.12]). Hence $\partial B(p_{j},r_{p})$ are locally uniformly contractible. Furthermore, by the stability of local uniform contractibility ([17, Section 16, Main Theorem]; see also [82, Section 5, Theorem], [42, Theorem 1.2]), we see that $\partial B(p,r_{p})$ is also locally uniformly contractible.
@@
Let $X\in\mathcal{G}(r,L)$. By Lemma 7.14 and the injectivity radius bound, every point of $X$ has a tiny ball of radius $r/100$. By the proof of Proposition 3.1 and the packing number bound, every tiny ball in $X$ of radius $r/100$ is $L$-doubling. In particular, the dimension of $X$ is bounded above by a constant depending only on $L$. Since $X$ is covered by at most $L$ tiny balls of radius $r/100$, this implies that for any given $\varepsilon>0$, the minimal number of $\varepsilon$-balls needed to cover $X$ is bounded above by a constant $N_{L}(\varepsilon)$ depending only on $L$ and $\varepsilon$. Therefore, the family $\mathcal{G}(r,L)$ is precompact in the Gromov–Hausdorff topology (cf. [63, Proposition 5.10]).
@@ 7.4. Generalization
On the other hand, the non-uniqueness of shortest paths typically results in the non-existence of a unique projection to a “$\sigma$-convex subset” (compare [78, Corollary 8.2.6, Proposition 8.4.8]). However, we have never used such a projection.
The advantage of dropping the uniqueness of shortest paths is that this weaker notion is preserved under the Gromov–Hausdorff convergence, which was essential in the previous subsection. Compare the following with [63, Example 4.3].
@@ 8. Problems
The first is the generalization of upper curvature bounds mentioned in Section 1.5. By modifying the BNPC condition in terms of comparison angles, one can define Busemann spaces with curvature $\leq\kappa$ for arbitrary $\kappa\in\mathbb{R}$ (see Remark 2.5).
@@
The next three problems are all related to the tangent cones of GNPC spaces, which also generalize previous results for GCBA spaces by Lytchak–Nagano [63]. Before discussing the problems, we briefly review the original results in [63] (cf. [77], [75], [15]). For analogous results for Alexandrov spaces with curvature bounded below, see [76], [74], [75], [81], [13]. We also refer to [61] for tangent cones in more general settings.
@@
We say that $p$ is a $k$-regular point if $T_{p}X$ is isometric to $\mathbb{R}^{k}$ (with Euclidean metric). The set $R^{k}$ of $k$-regular points in $X$ is dense and has full measure with respect to the $k$-dimensional Hausdorff measure in the $k$-dimensional part $X^{k}$ of $X$. Here, the $k$-dimensional part $X^{k}$ is the set of points in $X$ with $\dim T_{p}X=k$ (in the sense of the Hausdorff or topological dimension).
@@
We want to develop GNPC analogs of the above results. For simplicity, we ignore the dimensional inhomogeneity of a GNPC space and formulate very rough problems. Note that these are not rigorous conjectures.
The starting point would be the structure of the tangent cone (cf. [61]).
@@
Describe the structure of the tangent cone of a GNPC space. Define regular points of a GNPC space and show its almost everywhere existence.
Considering that the tangent space of a Finsler manifold is a strictly convex normed space, a regular point of a GNPC space should be defined as a point at which the tangent cone is isometric to a strictly convex normed space. Note that, unlike a Riemannian metric modeled on the standard Euclidean metric, a Finsler norm generally varies pointwise. However, if a Finsler manifold satisfies the BNPC condition, then the Finsler metric belongs to a special class called a Berward metric, as shown by Ivanov–Lytchak [51]. One of the properties of a Berwald metric is that the norm of the tangent space is all isometric. From our viewpoint that GNPC spaces is a generalization of Finsler manifolds with nonpositive flag curvature, one might ask the following.
@@
Does a GNPC space satisfy any kind of Berwardness? For example, is it possible to “decompose” a GNPC space so that in each piece the tangent cone is isometric to a fixed strictly convex normed space at almost every point?
Related problems and partial counterexamples can be found in [51, Sections 1, 6]. For example, one can glue two different strictly convex normed spaces at one point to obtain a GNPC space such that the tangent cones at “regular points” are not isometric. However, the authors do not know more complicated counterexamples to the above problem.
Note that, unlike Riemannian manifolds, Finsler manifolds do not generally have a “canonical” measure, but for (reversible) Berwald spaces the Hausdorff measure can be regarded as a canonical measure (see [73, Section 10.2]). Therefore it would be natural to consider the Hausdorff measure for our GNPC spaces.
Furthermore, based on the notion of a regular point, one could ask the following. Compare with Theorem 5.30 and Remark 5.31.
@@
Does a GNPC space admit any kind of Finsler structure almost everywhere? Does a regular point have a neighborhood almost isometric to an open subset of its tangent cone?
@@
Another problem we would like to mention here is the diffeomorphism version of Theorem 1.3 in the cocompact setting. As mentioned in [30, Section 6], this is open even in the CAT(00) case. See also [89] and [88] for related results.
@@
Is the universal cover of a closed smooth $4$-manifold with a locally CAT(00)/BNPC metric diffeomorphic to the standard $\mathbb{R}^{4}$?
Finally, we discuss Busemann spaces with lower curvature bounds. Recall that Theorems 1.1 and 1.2 were originally proved by Wu [96] for Alexandrov spaces with curvature bounded below. Reversing the inequality for comparison angles in Remark 2.5, one can also define Busemann spaces with curvature $\geq\kappa$, where $\kappa\in\mathbb{R}$ (cf. [44]; however, note that there is no equivalent definition in terms of the concavity of the distance function, similar to Definition 2.1, even for nonnegative curvature). Clearly, every Alexandrov space is a Busemann space with the same lower curvature bound. Therefore, it is natural to ask
@@
Note that Wu’s proof [96] (cf. [97]) relies heavily on the structure theorem for Alexandrov spaces by Perelman [79], [80], which seems very hard to generalize to Busemann spaces. We point out that Lytchak–Nagano’s proof [64] can also be applied to Alexandrov spaces, by using the gradient flow of the distance function (see [83]) instead of the geodesic contraction we frequently used. However, to the best of our knowledge, it is not even known whether a Busemann space with a lower curvature bound is locally contractible. We remark that there is work of Kell [53] on Busemann spaces with nonnegative curvature.